ANALYSIS OF THE MAGNETIC HYPERFINE STRUCTURE IN THE B^{3}Φ_{4} − X^{3}Φ_{4} BAND SYSTEM OF IRIDIUM MONOFLUORIDE, IrF

\underline{HAILING WANG}, XIUJUAN ZHUANG AND TIMOTHY C. STEIMLE , Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287; COLAN LINTON, Center for Lasers, Atomic and Molecular Sciences and Physics Department, University of New Brunswick, Fredericton, NB Canada E3B 5A3.

Recently the New Brunswick group^{a} reported on the detection and analysis of the B^{3}Φ_{4} − X^{3}Φ_{4} band system of IrF. The free-jet expansion conditions limited the spectral resolution to approximately 200 MHz, which was insufficient to fully resolve the ^{191}Ir(I=3/2), ^{193}Ir(I=3/2)and ^{19}F(I=1/2)magnetic hyperfine splitting. Here we report on the analysis of the same band system recorded under molecular beam conditions which resulted in a resolution of 40 MHz. A simple molecular orbital correlation diagram is used to rationalize the determined parameters of IrF and the isovalent molecules RhF ^{b} and CoF^{c}.

aA.G. Adam; A.D. Granger; L.E. Downie; D.W. Tokaryk and C. Linton Can.J. Phys. (accepted).

bR. Li, R.J. Jensen, W.J. Balfour, S.A. Shepard and A.G. Adam; J.Chem.Phys. 121 2591, 2004.

cJ.J. Harrison, J.M. Brown, M.A. Flory, P.M. Sheridan, S.K. McLamarrah and L.M. Ziurys J. Chem. Phys. 127 194308, 2007.